# On the distance between mean and geometric median in high dimensions

### v2

@@ 1 Introduction
The geometric median provides a robust alternative to the sample mean, mitigaging the impact of outliers. The improved robustness comes at the cost of a bias: in the infinite-data limit, the geometric median is generally distinct from the mean (i.e., expectation) of the data generating distribution, which in practice often is the learning target [^1] For special distributions like the Gaussian, geometric median and mean can coincide. . A method to balance robustness and bias is the median-of-means [8, 10] or “bucketing” approach [5].
In this paper, we take a step back and quantify the actual bias of the geometric median, focusing on high-dimensional distributions. We prove, when each distribution component only depends on a fixed number of neighboring components, that the maximum-norm distance between geometric median and mean vector is of order $\mathcal{O}(\tfrac{1}{p})$, where $p$ denotes the dimension. Furthermore, we show that this rate is sharp in general.
The results are discussed at the end of the paper. Proofs are deferred to the appendix.
@@ 2 Theoretical results
We consider random vectors $\mathbf{X}:=(X_{1},\ldots,X_{p})\in\mathbb{R}^{p}$ for $p\in\mathbb{N}$, where the sequence of component random variables $(X_{i})_{i\in\mathbb{N}}$ is from the class $\mathcal{D}\equiv\mathcal{D}(M,q,c,\sigma_{\min},C)$ introduced below. A geometric median of $\mathbf{X}$ is any vector $\mathbf{m}\in\mathbb{R}^{p}$ minimizing the expectation $\operatorname{E}\left[\,\|\mathbf{X}-\mathbf{m}\|\,\right]$ for the Euclidean norm $\|\cdot\|$. By properties of the class $\mathcal{D}$, there exists a unique geometric median (Lemma A.1), which we denote by $\mathbf{m}_{p}$. We then study $\|\mathbf{m}_{p}-\boldsymbol{\mu}\|$, where $\boldsymbol{\mu}:=\operatorname{E}\left[\,\mathbf{X}\,\right]$.
@@
Let $M\geq 0$ be an integer, and let $c,\sigma_{\min}>0$, $q\geq 2$ and $C\geq 1$ be reals. The class $\mathcal{D}(M,q,c,\sigma_{\min},C)$ is comprised of all sequences of absolutely continuous r.v. $(X_{i})_{i\in\mathbb{N}}$ that are $M$-dependent, with $\operatorname{E}\left[\,|X_{i}|^{q}\,\right]\leq C$, $\operatorname{E}\left[\,|X_{i}-\mu_{i}|^{q}\,\right]\leq C$, and $\operatorname{sd}\left[\,X_{i}\,\right]\geq\sigma_{\min}>0$, and for which the probability density of $X_{i}$ is bounded by $c<\infty$ for all $i\in\mathbb{N}$.
Recall that $M$-dependence requires independence of $(X_{i})_{i\leq k}$ and $(X_{i})_{i\geq k+M+1}$ for all $k\in\mathbb{N}$. Note that $|\mu_{i}|\leq C$ and $\operatorname{Var}\left[\,X_{i}\,\right]\leq C$ since $C\geq 1$.
We build on the following central identity, derived from $\nabla_{\mathbf{m}}\operatorname{E}\left[\,\|\mathbf{X}-\mathbf{m}\|\,\right]|_{\mathbf{m}=\mathbf{m}_{p}}=0$ by interchanging derivative with expectation, and then adding $\pm\boldsymbol{\mu}$:
$\mathbf{m}_{p}-\boldsymbol{\mu}=\operatorname{E}\left[\,\frac{1}{\|\mathbf{X}-\mathbf{m}_{p}\|}\,\right]^{-1}\operatorname{E}\left[\,\frac{\mathbf{X}-\boldsymbol{\mu}}{\|\mathbf{X}-\mathbf{m}_{p}\|}\,\right],$ (1)
where all expectations exist when $p$ is large enough (Lemma A.2).
The intuition is now that, by $M$-dependence, $X_{i}-\mu_{i}$ and $\|\mathbf{X}-\mathbf{m}_{p}\|$ are nearly independent for $p$ large, and thus $m_{p,i}-\mu_{i}\approx\operatorname{E}\left[\,X_{i}-\mu_{i}\,\right]\cdot 1=0$ by (1). Formalizing this approach requires controlling $\operatorname{E}\left[\,\|\mathbf{X}-\mathbf{m}_{p}\|^{-1}\,\right]$, or slight variations involving $\check{\mathbf{X}}$ introduced below, uniformly in the dimension $p$. To this end, we combine the two lemmas, where the first is an initial bound.
@@
When $(X_{i})_{i}\in\mathcal{D}(M,q,c,\sigma_{\min},C)$, then $\|\mathbf{m}_{p}-\boldsymbol{\mu}\|\leq\sqrt{Cp}$.
@@
Let $(Y_{i})_{i\in\mathbb{N}}\in\mathcal{D}(M,q,c,\sigma_{\min},C)$ with $q>2$ and $\mathbf{v}_{p}\in\mathbb{R}^{p}$ with $\|\mathbf{v}_{p}\|^{2}\leq C_{2}p$ for $p\in\mathbb{N}$ and some $C_{2}<\infty$. With $\mathbf{Y}:=(Y_{1},\ldots,Y_{p})$ we have
$\sup_{p\geq\lceil 4k\rceil(M+1)+1}\operatorname{E}\left[\,\left(\frac{\|\mathbf{Y}-\mathbf{v}_{p}\|^{2}}{\operatorname{E}\left[\,\|\mathbf{Y}-\mathbf{v}_{p}\|^{2}\,\right]}\right)^{-k}\,\right]=\beta(k,C_{2},\mathcal{D})<\infty,$ (2)
for any $k>0$ and constant $\beta$ depending only on $k,C_{2}$ and the $\mathcal{D}$-parameters.
Both Lemmas together with the Cauchy-Schwarz inequality applied to the second expectation in (1) lead to the following intermediate bound.
Lemma 2.4.
If $(X_{i})_{i}\in\mathcal{D}(M,q,c,\sigma_{\min},C)$ with $q>2$, then for a constant $B_{1}$ only depending on the $\mathcal{D}$-parameters, $\sup_{p\in\mathbb{N};\;1\leq i\leq p}|\mu_{i}-m_{p,i}|\leq B_{1}$.
We can now derive the main results, which refine the Cauchy-Schwarz bound in the proof of Lemma 2.4. More precisely, we show that $\operatorname{E}\left[\,(X_{i}-\mu_{i})/\|\mathbf{X}-\mathbf{m}_{p}\|\,\right]$ is close to zero by inserting $0=\operatorname{E}\left[\,(X_{i}-\mu)/\|\check{\mathbf{X}}-\check{\mathbf{m}}_{p}\|\,\right]$, where $\check{\mathbf{X}}$ is missing all components with indices in $i-M,\ldots,i+M$. Formally, using the exclusion set $\mathcal{E}(i):=\mathbb{N}\cap\{i-M,\ldots,i+M\}$, we define $\check{\mathbf{v}}:=(v_{i})_{i\in\{1,\ldots,p\}\setminus\mathcal{E}(i)}\in\mathbb{R}^{|\{1,\ldots,p\}\setminus\mathcal{E}(i)|}$ for any vector $\mathbf{v}=(v_{i})_{i=1}^{p}\in\mathbb{R}^{p}$. Thus, subtracting zero from the right hand side of (1), a bit of algebra yields that for $p\geq i+M$,
$m_{p,i}-\mu_{i}=\operatorname{E}\left[\,\frac{1}{\|\mathbf{X}-\mathbf{m}_{p}\|}\,\right]^{-1}\operatorname{E}\left[\,\frac{-(X_{i}-\mu_{i})\sum_{j\in\mathcal{E}(i)}(X_{j}-m_{p,j})^{2}}{\|\mathbf{X}-\mathbf{m}_{p}\|\|\check{\mathbf{X}}-\check{\mathbf{m}}_{p}\|(\|\mathbf{X}-\mathbf{m}_{p}\|+\|\check{\mathbf{X}}-\check{\mathbf{m}}_{p}\|)}\,\right].$ (3)
The first expectation is $\mathcal{O}(\sqrt{p})$ while the second is $\mathcal{O}(\tfrac{1}{p^{3/2}})$, leading to
@@
Let $(X_{i})_{i}\in\mathcal{D}(M,q,c,\sigma_{\min},C)$ with $q\geq 3$. Then, with $B_{3}$ depending only on the $\mathcal{D}$-parameters, $|m_{p,i}-\mu_{i}|\leq\frac{B_{3}}{p}$ for all $1\leq i\leq p\in\mathbb{N}$.
In particular, Theorem 2.5 implies $\|\mathbf{m}_{p}-\boldsymbol{\mu}\|\leq\tfrac{B_{3}}{\sqrt{p}}$.
By $\|\mathbf{m}_{p}-\boldsymbol{\mu}\|\to 0$, we can identify the limit in probability of the (suitably scaled) expectation arguments in (3). A uniform integrability argument (requiring $q>3$) transfers the convergence to the expectations themselves and yields
@@
Let $(X_{i})_{i}\in\mathcal{D}(M,q,c,\sigma_{\min},C)$ with $q>3$. Fix $i\in\mathbb{N}$. If $\tfrac{1}{p}\sum_{i=1}^{p}\operatorname{Var}\left[\,X_{i}\,\right]\to\bar{\sigma}^{2}\in\mathbb{R}$, then
$m_{p,i}=\mu_{i}-\frac{1}{2\bar{\sigma}^{2}p}\sum_{j\in\mathcal{E}(i)}\operatorname{E}\left[\,(X_{i}-\mu_{i})(X_{j}-\mu_{j})^{2}\,\right]+o\left(\frac{1}{p}\right).$ (4)
Note that while $\sigma_{\min}^{2}\leq\tfrac{1}{p}\sum_{i=1}^{p}\operatorname{Var}\left[\,X_{i}\,\right]\leq C$ always has a convergent subsequence, there may be multiple accumulation points.
@@ 3 Simulations
We simulate $\hat{\mathbf{m}}_{p}$ for a range of distributions to validate Theorems 2.5 and 2.6. The simulation details are discussed at the end of this section.
@@
$\left|\hat{m}_{p,1}-\mu_{1}+\frac{1}{2\bar{\sigma}^{2}p}\sum_{j\in\mathcal{E}(1)}\operatorname{E}\left[\,(X_{1}-\mu_{1})(X_{j}-\mu_{j})^{2}\,\right]\right|,$ (5)

### v3

@@ 1 Introduction
The geometric median provides a robust alternative to the sample mean, mitigaging the impact of outliers. The improved robustness comes at the cost of a bias: in the infinite-data limit, the geometric median is generally distinct from the mean (i.e., expectation) of the data generating distribution, which often is the learning target in practice [^1] For special distributions like the Gaussian, geometric median and mean can coincide. . A method to balance robustness and bias is the median-of-means [10, 12] or “bucketing” approach [7].
In this paper, we take a step back and quantify the actual bias of the geometric median, focusing on high-dimensional distributions. We prove, when each distribution component only depends on a fixed number of neighboring components, that the maximum-norm distance between geometric median and mean vector is of order $\mathcal{O}(\tfrac{1}{p})$, where $p$ denotes the dimension. Furthermore, we derive an explicit expansion of the geometric median components when $p$ tends to infinity, in particular showing $\mathcal{O}(\tfrac{1}{p})$ is sharp in general.
We discuss connections to the literature and implications of our results at the end of the paper. Proofs are deferred to the appendix.
2 Background and notation
Let $\mathbf{X}:=(X_{1},\ldots,X_{p})\in\mathbb{R}^{p}$ be a $p$-variate random vector, and $\mathbf{x}_{1},\ldots,\mathbf{x}_{n}$ be independent samples from $\mathbf{X}$. We denote the Euclidean norm by $\|\cdot\|$. Consider the two objective functions
$D(\mathbf{m}):=\operatorname{E}\left[\,\|\mathbf{X}-\mathbf{m}\|\,\right]$ and $\hat{D}(\mathbf{m}):=\frac{1}{n}\sum_{i=1}^{n}\|\mathbf{x}_{i}-\mathbf{m}\|.$
Any vector $\mathbf{m}\in\mathbb{R}^{p}$ minimizing $D$ or $\hat{D}$ is called population or sample geometric median of $\mathbf{X}$, respectively. A minimizer exists under mild conditions; in fact, by replacing $D(\mathbf{m})$ with the equivalent objective $\operatorname{E}\left[\,\|\mathbf{X}-\mathbf{m}\|-\|\mathbf{X}\|\,\right]$, moment conditions on $\mathbf{X}$ can be avoided. When the minimizer is unique, we write $\mathbf{m}$ for the population and $\hat{\mathbf{m}}$ for the sample geometric median. By translation invariance of $D$, one can shift $\mathbf{m}$ to the origin without loss of generality, however we keep $\mathbf{m}$ as a variable to make the distance between median and mean recognizable in the formulas.
The geometric median $\hat{\mathbf{m}}$ has light tails even under minimal moment assumptions on $\mathbf{X}$ [12]. This property makes it popular in heavy-tailed location testing where it can help outperform alternatives like Hotelling’s $T^{2}$ [6, 5], and more generally in robust statistics and machine learning [7].
We write $\boldsymbol{\mu}:=\operatorname{E}\left[\,\mathbf{X}\,\right]$ for the population mean and $\operatorname{sd}\left[\,X_{i}\,\right]$ and $\operatorname{Var}\left[\,X_{i}\,\right]$ for the standard deviation and variance of component $X_{i}$. For $\mathbf{v}\in\mathbb{R}^{p}$, we let $\|\mathbf{v}\|_{\infty}:=\max(|v_{1}|,\ldots,|v_{p}|)$. We use the Landau symbols $\mathcal{O}(\cdot)$ and $o(\cdot)$.
@@ 3 Theoretical results
This section examines the distance between mean and geometric median and obtains the results highlighted in Table 1. We begin with our distributional assumptions.
Recall that a sequence of random variables $(X_{i})_{i\in\mathbb{N}}$ is $M$-dependent if $(X_{i})_{i=1}^{k}$ and $(X_{i})_{i=k+M+1}^{\infty}$ are independent for all $k\in\mathbb{N}$. $M$-dependence of finite-dimensional random vectors $\mathbf{X}=(X_{i})_{1\leq i\leq p}$ is defined analogously under restriction to subvectors of $\mathbf{X}$.
@@
Let $M\geq 0$ be an integer, and let $c,\sigma_{\min}>0$, $q\geq 2$, and $C\geq 1$ be reals. The distribution of an $M$-dependent random vector $\mathbf{X}=(X_{i})_{1\leq i\leq p}$ or an $M$-dependent sequence $(X_{i})_{i\in\mathbb{N}}$ belongs to the class $\mathcal{D}(M,q,c,\sigma_{\min},C)$ if all components $X_{i}$ admit a Lebesgue density bounded by $c$, with $\operatorname{E}\left[\,|X_{i}|^{q}\,\right]\leq C$ and $\operatorname{sd}\left[\,X_{i}\,\right]\geq\sigma_{\min}>0$.
Note that the definition implies that $|\mu_{i}|\leq C$ and $\operatorname{Var}\left[\,X_{i}\,\right]\leq C$ as $C\geq 1$.
Result Integrability Reference
$\|\mathbf{m}-\boldsymbol{\mu}\|_{\infty}\leq\text{const}$ $q>2$ Lemma 3.3
$\|\mathbf{m}-\boldsymbol{\mu}\|_{\infty}=\mathcal{O}(\tfrac{1}{p})$ $q\geq 3$ Theorem 3.4
$m_{p,i}=\mu_{i}-\tfrac{\alpha_{i}}{p}+o(\tfrac{1}{p})$ $q>3$ Theorem 3.5
Table 1: Theoretical results with required integrability $q$.
Let the $p$-variate random vector $\mathbf{X}$ belong to the class $\mathcal{D}\equiv\mathcal{D}(M,q,c,\sigma_{\min},C)$. Then, $\mathbf{X}$ is guaranteed to have a unique population geometric median $\mathbf{m}$ when $p\geq M+2$ (Lemma A.1). We build on the following identity for $\mathbf{m}$, which is obtained by setting derivatives in the geometric median objective to zero:
$\mathbf{m}-\boldsymbol{\mu}=\operatorname{E}\left[\,\frac{1}{\|\mathbf{X}-\mathbf{m}\|}\,\right]^{-1}\operatorname{E}\left[\,\frac{\mathbf{X}-\boldsymbol{\mu}}{\|\mathbf{X}-\mathbf{m}\|}\,\right].$ (1)
Equation (1) can be proven rigorously when $p$ is large enough (Lemma A.2).
The guiding intuition for why $\|\mathbf{m}-\boldsymbol{\mu}\|$ should be small is that, by $M$-dependence, $X_{i}-\mu_{i}$ and $\|\mathbf{X}-\mathbf{m}\|$ are nearly independent for large $p$. Were they actually independent, the second expectation in (1) would factorize and imply $\mathbf{m}-\boldsymbol{\mu}=\mathbf{0}$, since $\operatorname{E}\left[\,X_{i}-\mu_{i}\,\right]=0$.
Formalizing this approach requires controlling $\operatorname{E}\left[\,\|\mathbf{X}-\mathbf{m}\|^{-1}\,\right]$, or slight variations introduced later, uniformly in the dimension $p$. To deal with $\mathbf{m}$ on the right hand side of (1), we use the standard a-priori bound $\|\mathbf{m}-\boldsymbol{\mu}\|^{2}\leq C\,p$ (see Lemma A.3 or Minsker and Strawn 13).
@@
Let $\mathbf{Y}:=(Y_{1},\ldots,Y_{p})\in\mathcal{D}(M,q,c,\sigma_{\min},C)$ with $q>2$, and let $\mathbf{v}\in\mathbb{R}^{p}$ with $\|\mathbf{v}\|^{2}\leq C_{2}\,p$ for some $C_{2}<\infty$. When $p\geq\lceil 4k\rceil(M+1)+1$,
$\operatorname{E}\left[\,\left(\frac{1}{\|\mathbf{Y}-\mathbf{v}\|^{2}}\right)^{k}\,\right]\leq\frac{\beta(k,C_{2},\mathcal{D})}{\operatorname{E}\left[\,\|\mathbf{Y}-\mathbf{v}\|^{2}\,\right]^{k}}<\infty$ (2)
for any $k>0$ and a constant $\beta$ depending only on $k,C_{2}$ and the $\mathcal{D}$-parameters.
Applying the Cauchy-Schwarz inequality to the second expectation in (1) and inserting Lemma 3.2 yields the following very useful intermediate bound.
@@
Let the $p$-variate $\mathbf{X}$ belong to $\mathcal{D}(M,q,c,\sigma_{\min},C)$ with $q>2$. Then, $\|\mathbf{m}-\boldsymbol{\mu}\|_{\infty}\leq B_{1}$ for a constant $B_{1}$ depending only on the $\mathcal{D}$-parameters.
For deriving the main results, we measure the distance between each right hand side component of equation (1) and zero in a special way. Precisely, let $i\leq p$ be fixed and define $\check{\mathbf{X}}:=(X_{j})_{\{1\leq j\leq p\,:\,|j-i|>M\}}$. With $\check{\mathbf{m}}$ defined similarly, $\operatorname{E}\left[\,\nicefrac{{(X_{i}-\mu_{i})}}{{\|\check{\mathbf{X}}-\check{\mathbf{m}}\|}}\,\right]=0$ by $M$-dependence. Subtracting this zero from equation (1), we obtain after a bit of algebra that
$m_{i}-\mu_{i}=\operatorname{E}\left[\,\frac{1}{\|\mathbf{X}-\mathbf{m}\|}\,\right]^{-1}\operatorname{E}\left[\,\frac{-(X_{i}-\mu_{i})\sum_{|j-i|\leq M}(X_{j}-m_{j})^{2}}{\|\mathbf{X}-\mathbf{m}\|\|\check{\mathbf{X}}-\check{\mathbf{m}}\|(\|\mathbf{X}-\mathbf{m}\|+\|\check{\mathbf{X}}-\check{\mathbf{m}}\|)}\,\right],$ (3)
as proven for $p$ large enough in Lemma A.4. The first expectation is $\mathcal{O}(\sqrt{p})$ while the second is $\mathcal{O}(p^{-3/2})$, leading to the first main result:
@@
Let the $p$-variate $\mathbf{X}$ belong to $\mathcal{D}(M,q,c,\sigma_{\min},C)$ with $q\geq 3$. Then, $\|\mathbf{m}-\boldsymbol{\mu}\|_{\infty}\leq\tfrac{B_{3}}{p}$ for $B_{3}$ depending only on the $\mathcal{D}$-parameters.
Next, we consider random vectors $\mathbf{X}_{p}\in\mathbb{R}^{p}$ and their medians $\mathbf{m}_{p}$ as $p$ tends to infinity. We establish an asymptotic expansion $m_{p,i}=\mu_{i}-\tfrac{\alpha_{i}}{p}+o(\tfrac{1}{p})$ under mild assumptions. This shows that the rate obtained in Theorem 3.4 is generally sharp, and through explicit formulas for $\alpha_{i}$ we gain insight into which distributional properties of $\mathbf{X}$ influence $m_{p,i}-\mu_{i}$ most strongly.
We prove that $p\cdot(m_{p,i}-\mu_{i})$ converges to a constant $\alpha_{i}$ as $p\to\infty$. Precisely, the random variables inside the expectations on the right hand side of the scaled equation (3) admit a limit in probability, since $m_{p,i}\to\mu_{i}$ by Theorem 3.4. We transfer this convergence in probability to convergence of the expectations via a uniform integrability argument, requiring $q>3$.
@@
Let $(X_{i})_{i\in\mathbb{N}}\in\mathcal{D}(M,q,c,\sigma_{\min},C)$ with $q>3$ and assume $\tfrac{1}{p}\sum_{i=1}^{p}\operatorname{Var}\left[\,X_{i}\,\right]\to\bar{\sigma}^{2}$. Fix $i\in\mathbb{N}$. Then, $m_{p,i}=\mu_{i}-\tfrac{\alpha_{i}}{p}+o(\tfrac{1}{p})$, where
$\alpha_{i}:=\frac{1}{2\bar{\sigma}^{2}}\sum_{|j-i|\leq M}\operatorname{E}\left[\,(X_{i}-\mu_{i})(X_{j}-\mu_{j})^{2}\,\right].$
The expansion coeffients $\alpha_{i}$ can be computed analytically in many cases. When $M=0$, i.e., $(X_{i})_{i}$ are independent, the sum reduces to the skewness of $X_{i}$. Note that while $\sigma_{\min}^{2}\leq\tfrac{1}{p}\sum_{i=1}^{p}\operatorname{Var}\left[\,X_{i}\,\right]\leq C$ always has a convergent subsequence, there may be multiple accumulation points. In this case, the expansion is satisfied on each convergent subsequence with individual expansion coefficients $\alpha_{i}$.
@@ 4 Simulations
We simulate $\hat{\mathbf{m}}_{p}$ for a range of distributions to validate Theorems 3.4 and 3.5. The simulation details are discussed at the end of this section.
@@
$\big|\hat{m}_{p,1}-\mu_{1}+\frac{1}{2\bar{\sigma}^{2}p}\sum_{j=1}^{M}\operatorname{E}\left[\,(X_{1}-\mu_{1})(X_{j}-\mu_{j})^{2}\,\right]\big|,$ (4)